Hey folks,
Got a problem here. The area of a rectangle is 308 square feet, and is 8 feet longer than the width. My teacher has supplied the answer: 14 feet by 22 feet, but I can't figure out how to arrive at those numbers. Help!!!
Mike

Answer by checkley77(12844)

AREA=LW
Given: Area=308 & Length is the Width plus 8 (W+8).
L=W+8
Replace L by (W+8) & solve for W.
308=(W+8)W
308=W^2+8W
W^2+8W-308=0
(W+22)(W-14)=0
W-14=0
W=14 ANS. FOR THE WIDTH.
14=W+8
L=14+8=22 ANS. FOR THE LENGTH.
PROOF:
308=22*14
308=308